OpenStudy (anonymous):

Solve the equation 7/3 - 1/x-1 = -3/3x-3 . Use complete sentences to explain how you solved this equation.

OpenStudy (cwrw238):

you can first simplify it a bit by making the RHS = -1/ x - 1 and that gives 7/3 = 0

hero (hero):

@cwrw238 is correct. You would that, however, you would have to conclude that \(\frac{7}{3} \ne 0\)

OpenStudy (anonymous):

thanks, that was really helpful

OpenStudy (anonymous):

can you explain it in a little more detail please

hero (hero):

Yes, I can explain all the steps. 1. You want to make sure all of the denominators are factored: \(\large\frac{7}{3} - \frac{1}{x-1} = -\frac{3}{3(x-1)}\) 2. Next cancel 3/3 on the right side to get: \(\large\frac{7}{3} - \frac{1}{x-1} = -\frac{1}{x-1}\) 3. Next add 1/(x-1) to both sides to get: \(\large\frac{7}{3} = -\frac{1}{x-1} + \frac{1}{x-1}\) 4. Next combine fractions to get: \(\large\frac{7}{3} = \frac{-1+1}{x-1} \) 5. Next simplify to get: \(\large\frac{7}{3} = \frac{0}{x-1} \) 6. You know that any fraction with zero in the numerator is zero, so: \(\large\frac{7}{3} = 0 \) 7. You know that 7/3 = 0 is false, so : \(\frac{7}{3} \ne 0\) 8. Therefore No Solution

OpenStudy (anonymous):

Thanks your a life saver